OpenStudy (anonymous):

Angle 1 and angle 2 are supplementary angles. the measure of angle 1 is 4y+7, and the measure of angle 2 is 9y+4. Find the measure of angle 2.

OpenStudy (anonymous):

Do u want angle 2 ?

OpenStudy (anonymous):

yes, can you help?

OpenStudy (anonymous):

Yep ! If equation of 1 and 2 we have ,we can have an equation from both of them that s 180 OK ?

OpenStudy (anonymous):

right because it is a supplementary angle, so it would have to add up to 180, but I don't understand how to get the angle measurements.

OpenStudy (anonymous):

No problem :) Step by step :D Now can u answer this : 4y+7+9y+4=?

OpenStudy (anonymous):

no, not without y

OpenStudy (anonymous):

ooh wait it would equal 180

OpenStudy (anonymous):

So smart ! Yea !

OpenStudy (anonymous):

If u do that have an equal from 180 :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Not thing ... Just answer to this if u can : 4y+7+9y+4=?

OpenStudy (anonymous):

the equation 4y+7+9y+4=180 But how do you figure from there?

OpenStudy (anonymous):

oooh wait what if I combine the commons together like 4y and 9y

OpenStudy (anonymous):

Exacly :)

OpenStudy (anonymous):

ok so now we would have 13y+11, right?

OpenStudy (anonymous):

Exacly !

OpenStudy (anonymous):

ok :D Now I'm lost, I don't know what to do from here.

OpenStudy (anonymous):

Look we have : 13y+11=180 right ? So now if u find y can find the angles :) Just find y now :)

OpenStudy (anonymous):

ok, so if we have 13y+11=180 then I guess I would have to get rid of the 11. To do that I have to subtract it away and subtract it to the other side to even it out. So after that we would have 13y=169.

OpenStudy (anonymous):

Now I can divide 169 by 13 to get y. So when I divide 169 by 13 I get 13, so would that be my answer to the y in the equations to the angles?

OpenStudy (anonymous):

Exelent :) Now we have : angle 2 = 9y+4 Find it :)

OpenStudy (anonymous):

Well if 9y+4 is equal to the measurement to angle 2 then to find this I just plug 13 in for y and I now have 9*13+4. 9*13 is 117. Now I have to add the 4 to 117, and that would be 121, so the angle measurement to 2 would be 121?

OpenStudy (anonymous):

Exelent !:) Your answer is realy right and you re realy smart too :)

OpenStudy (anonymous):

Thnx a lot I owe you this :D

OpenStudy (anonymous):

Thank u :) I do not think ... You answer to your self :)

OpenStudy (anonymous):

really? because it didn't feel that way at all

OpenStudy (anonymous):

But I feel at all :) You never need my help :)

OpenStudy (anonymous):

haha I might on the next one I post :P

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

Here is another one that I am stuck on, GI bisects <DGH so that m <DGI is x-3 and m <IGH is 2x-13 find the value of X.

OpenStudy (anonymous):

Oh excuse me but I should to go but u can say me latter or help from another ... I m realy sorry ... I think @PFEH.1999 can help u ... Bye and nice to talk u ..:)

OpenStudy (anonymous):

Thank you again :D have a good day